You are working on an IPO for a tech company. It is expected that the company...

  1. You are working on an IPO for a tech company. It is expected that the company will have earnings of $10 per share next year. You are required to estimate its value. You have the following information about its peers. What would be a reasonable estimate of the value of this company? Show all calculations.

Peers

Earnings per share

Stock Price

A

$10

$500

B

$15

$800

C

$20

$1000

D

$5

$250

E

$1

$100
Answer #1

Average Price/Earnings=(500/10+800/15+1000/20+250/5+100/1)/5=60.66666667

Hence, Value=Price/Earnings*Earnings=60.66666667*10=606.6666667